# An Actionable Framework for Assessing Bias and Fairness in Large Language Model Use Cases

### v2

@@ 3.1.3.2 Counterfactual Sentiment Bias
Counterfactual sentiment metrics measure the sentiment consistency across counterfactually generated pairs of output. To achieve this, these metrics leverage a pre-trained sentiment classifier $Sm:\mathcal{Y}\xrightarrow[]{}[0,1]$. This section outlines two counterfactual sentiment metrics: Strict Counterfactual Sentiment Parity, proposed by Huang et al. (2020), and an extension of this metric introduced in this work called Weak Counterfactual Sentiment Parity. [^19] Huang et al. (2020) use Google Cloud sentiment API and a BERT-based sentiment classifier. Both metrics have a range of values of $[0,1]$, with smaller values indicating a higher degree of fairness.
@@
This study presents WCSP, defined as the difference in predicted sentiment rates by a sentiment classifier applied to counterfactually generated LLM output pairs. Given a threshold $\tau$ for binarizing sentiment scores, this metric is introduced as follows:
@@
3.2 Fairness Metrics for Classification Use Cases
It is well-established that classification models can produce unfair outcomes for certain protected attribute groups (Saleiro et al., 2018; Bellamy et al., 2018; Weerts et al., 2023; Feldman et al., 2014; Hardt et al., 2016; Mehrabi et al., 2019). Let a classification LLM use case be defined as an LLM tasked with classification, denoted as $\mathcal{M}^{(c)}$, and a population of prompts $\mathcal{P}_{X}$. Note that the metrics introduced in this section are confined to binary classification use cases, where $\mathcal{M}^{(c)}:\mathcal{X}\xrightarrow{}\{0,1\},$ with the understanding that evaluating fairness for multiclass classification is a straightforward extension from the binary case (Rouzot et al., 2023).
For the remainder of this section, assume each prompt in $\mathcal{P}_{X}$ for a given classification LLM use case corresponds to a protected attribute group. Under this assumption, traditional machine learning fairness metrics (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023) may be applied (Czarnowska et al., 2021). Accordingly, these metrics are defined on binary predictions $\hat{Y}_{1},...,\hat{Y}_{N}$, generated from a sample of prompts $X_{1},...,X_{N}\in\mathcal{P}_{X}$, with some metrics also incorporating corresponding ground truth values $Y_{1},...,Y_{N}$. These metrics effectively assess group fairness (see Section 2.2) between two protected attribute groups $G^{\prime}$ and $G^{\prime\prime}$, with choice of statistical outcome measure $B$ depending on stakeholder values (e.g. the relative cost of false negatives vs. false positives).
This section distinguishes between representation fairness metrics, calculated using only predictions, and error-based fairness metrics, calculated using both predictions and ground truth values. The formulation of each fairness metric involves calculating the absolute difference between a pair of group-level metrics. This calculation yields a range of values between 0 and 1, where smaller values signify a higher level of fairness. [^20] Note that while this work considers group fairness in terms of differences between protected attribute groups, as in Gallegos et al. (2023); Bellamy et al. (2018), analogous metrics can be calculated as ratios (Saleiro et al., 2018).
@@ 3.2.1 Representation Fairness Metrics for Binary Classification
Representation fairness metrics aim to determine whether protected attribute groups are adequately represented in the positive predictions generated by classifier. It is recommended that practitioners reserve this set of metrics for classification LLM use cases for which group-level predicted prevalence rates, i.e. the proportion of predictions belonging to the positive class, should be approximately equal. [^21] For example, when predicting which applicants are qualified for a job it may be desirable to target equal rates of positive predictions for males and females. However, when predicting certain diseases, this may not be a desirable model behavior. A single representation fairness metric, known as Demographic Parity, is defined below.
@@ 3.2.2 Error-Based Fairness Metrics for Binary Classification
Error-based fairness metrics aim to determine whether disparities in model performance exist across protected attribute groups. For error-based fairness, two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, are presented. Following Saleiro et al. (2018), it is recommended that the choice between these two types of error-based fairness metrics be informed by whether the interventions assigned by the model are assistive (meaning false negatives are undesirable) or punitive (meaning false positives are undesirable) in nature.
@@ 3.2.3 Multiclass Fairness Metrics
For multiclass classifiers, the framework proposed here follows the fairness guidelines provided by Rouzot et al. (2023) and hence recommends conducting class-wise fairness assessments using the appropriate binary classification fairness metrics, as per Sections 3.2.1, 3.2.2, on each of the ‘sensitive’ classes. In particular, Rouzot et al. (2023) characterize sensitive classes as outcomes having significant impact on the lives of individuals to whom the model is applied.
3.3 Fairness Metrics for Recommendation Use Cases
@@
This section outlines a set of fairness metrics for recommendation LLM use cases, as proposed by Zhang et al. (2023). To ensure consistency with the metrics discussed in Section 3.1.3, this section presents modified versions of these metrics to be pairwise in nature, rather than attribute-wise. Given two protected attribute groups $G^{\prime},G^{\prime\prime},$ and an LLM use case $(\mathcal{M}^{(R)},\mathcal{P}_{X})$, these metrics assess similarity in counterfactually generated recommendation lists. Below, each metric is defined according to responses generated from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ which are drawn from $\mathcal{P}_{X|\mathcal{A}}$. In particular, three metrics are presented: Jaccard Similarity, Search Result Page Misinformation Score at K, and Pairwise Ranking Accuracy Gap at K. Each of these metrics ranges in value from 0 to 1, with larger values indicating a greater degree of fairness.
@@
Below, a pairwise version of Jaccard-K is introduced. This metric calculates the average Jaccard Similarity (Han et al., 2011)—the ratio of the intersection cardinality to the union cardinality—among pairs of counterfactually generated recommendation lists. Formally, this metric is computed as follows:
@@
Adapted from Tomlein et al. (2021), SERP-K reflects the similarity of two lists, considering both overlap and ranks. A modified version of SERP-K, adapted for pairwise application, is introduced as follows:
@@
where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$, $v$ is a recommendation from $\hat{R}_{i}^{\prime}$, and $rank(v,\hat{R}_{i}^{\prime})$ denotes the rank of $v$ in $\hat{R}_{i}^{\prime}$. Note that the use of $\min(\cdot,\cdot)$ is included to achieve symmetry.
@@
Adapted from Beutel et al. (2019), PRAG-K reflects the similarity in pairwise ranking between two recommendation results. A pairwise version of PRAG-K is presented as follows:
@@
where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$, $v_{1},v_{2}$ are recommendations from $\hat{R}_{i}^{\prime}$, and $rank(v,\hat{R}_{i})$ denotes the rank of $v$ in $\hat{R}_{i}$. As with SERP-K, $\min(\cdot,\cdot)$ is used to achieve symmetry.
Table 2: Bias and Fairness Evaluation Metrics

### v3

@@ 3.1.3.2 Counterfactual Sentiment Bias
Counterfactual sentiment metrics measure the sentiment consistency across counterfactually generated pairs of output. To achieve this, these metrics leverage a pre-trained sentiment classifier $Sm:\mathcal{Y}\xrightarrow[]{}[0,1]$. We outline two counterfactual sentiment metrics: Strict Counterfactual Sentiment Parity, proposed by Huang et al. (2020), and an extension of this metric called Weak Counterfactual Sentiment Parity. [^23] Huang et al. (2020) use Google Cloud sentiment API and a BERT-based sentiment classifier. Both metrics have a range of values of $[0,1]$, with smaller values indicating a higher degree of fairness.
@@
We introduce WCSP, defined as the difference in predicted sentiment rates by a sentiment classifier applied to counterfactually generated LLM output pairs. Given a threshold $\tau$ for binarizing sentiment scores, this metric is introduced as follows:
@@
3.2 Metrics for Classification Use Cases
It is well-established that classification models can produce unfair outcomes for certain protected attribute groups (Saleiro et al., 2018; Bellamy et al., 2018; Weerts et al., 2023; Feldman et al., 2014; Hardt et al., 2016; Mehrabi et al., 2019). Let a classification LLM use case be defined as an LLM tasked with classification, denoted as $\mathcal{M}^{(c)}$, and a population of prompts $\mathcal{P}_{X}$. Here, we present metrics for binary classification use cases, where $\mathcal{M}^{(c)}:\mathcal{X}\xrightarrow{}\{0,1\},$ noting that evaluating fairness for multiclass classification is a straightforward extension from the binary case (Rouzot et al., 2023).
For the remainder of this section, we assume each prompt in $\mathcal{P}_{X}$ for a given classification LLM use case corresponds to a protected attribute group. Under this assumption, traditional machine learning fairness metrics (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023) may be applied (Czarnowska et al., 2021). Accordingly, we define these metrics on binary predictions $\hat{Y}_{1},...,\hat{Y}_{N}$, generated from a sample of prompts $X_{1},...,X_{N}\in\mathcal{P}_{X}$, with some metrics also incorporating corresponding ground truth values $Y_{1},...,Y_{N}$. These metrics effectively assess group fairness (see Section 2.2) between two protected attribute groups $G^{\prime}$ and $G^{\prime\prime}$, with choice of statistical outcome measure $B$ depending on stakeholder values (e.g. the relative cost of false negatives vs. false positives).
We distinguish between representation fairness metrics, calculated using only predictions, and error-based fairness metrics, calculated using both predictions and ground truth values. Each fairness metric measures the absolute difference between a pair of group-level metrics. This calculation yields a range of values between 0 and 1, where smaller values indicate a higher level of fairness. [^24] Note that while we consider group fairness in terms of differences between protected attribute groups, as in Gallegos et al. (2023); Bellamy et al. (2018), analogous metrics can be calculated as ratios (Saleiro et al., 2018).
@@ 3.2.1 Representation Fairness Metrics for Binary Classification
Representation fairness metrics aim to determine whether protected attribute groups are adequately represented in the positive predictions generated by a classifier. We recommend that practitioners reserve this set of metrics for classification LLM use cases for which group-level predicted prevalence rates, i.e. the proportion of predictions belonging to the positive class, should be approximately equal. [^25] For example, when predicting which applicants are qualified for a job it may be desirable to target equal rates of positive predictions for males and females. However, when predicting certain diseases, for instance, this may not be a desirable model behavior. In our framework, we include a single representation fairness metric, Demographic Parity, which measures the absolute difference in group-level predicted prevalence rates.
@@ 3.2.2 Error-Based Fairness Metrics for Binary Classification
Error-based fairness metrics aim to determine whether disparities in model performance exist across protected attribute groups. To address error-based fairness, we include two metrics focused on false negatives, False Negative Rate Difference and False Omission Rate Difference, and two metrics focused on false positives, False Positive Rate Difference and False Discovery Rate Difference, in our framework. Following Saleiro et al. (2018), we recommend that practitioners assess disparities in false negatives (positives) across groups for use cases assigning assistive (punitive) interventions. These metrics are defined below.
@@ 3.2.3 Multiclass Fairness Metrics
For multiclass classifiers, we follow the fairness guidelines provided by Rouzot et al. (2023). Hence, we recommend conducting class-wise, one-vs-rest, fairness assessments using the appropriate binary classification fairness metrics, as per Sections 3.2.1, 3.2.2, on each of the ‘sensitive’ classes. In particular, Rouzot et al. (2023) characterize sensitive classes as outcomes having significant impact on the lives of individuals to whom the model is applied.
3.3 Metrics for Recommendation Use Cases
@@
We outline a set of fairness metrics for recommendation LLM use cases, as proposed by Zhang et al. (2023). To maintain consistency with the metrics discussed in Section 3.1.3, we present modified versions of these metrics to be pairwise in nature, rather than attribute-wise. Given two protected attribute groups $G^{\prime},G^{\prime\prime},$ and an LLM use case $(\mathcal{M}^{(R)},\mathcal{P}_{X})$, these metrics assess similarity in counterfactually generated recommendation lists. Below, we define each metric according to responses generated from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ which are drawn from $\mathcal{P}_{X|\mathcal{A}}$. In particular, three metrics are presented: Jaccard Similarity, Search Result Page Misinformation Score at K, and Pairwise Ranking Accuracy Gap at K. Each of these metrics ranges in value from 0 to 1, with larger values indicating a greater degree of fairness.
@@
We present a pairwise version of Jaccard-K. This metric calculates the average Jaccard Similarity (Han et al., 2011)—the ratio of the intersection cardinality to the union cardinality—among pairs of counterfactually generated recommendation lists. Formally, this metric is computed as follows:
@@
Adapted from Tomlein et al. (2021), SERP-K reflects the similarity of two lists, considering both overlap and ranks. We define a modified version of SERP-K, adapted for pairwise application, is introduced as follows:
@@
where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$, $v$ is a recommendation from $\hat{R}_{i}^{\prime}$, and $rank(v,\hat{R}_{i}^{\prime})$ denotes the rank of $v$ in $\hat{R}_{i}^{\prime}$. Note that we use $\min(\cdot,\cdot)$ to achieve symmetry.
@@
Adapted from Beutel et al. (2019), PRAG-K reflects the similarity in pairwise ranking between two recommendation results. We define a pairwise version of PRAG-K as follows:
@@
where $\hat{R}_{i}^{\prime},\hat{R}_{i}^{\prime\prime}$ respectively denote the generated lists of recommendations by $\mathcal{M}(X;\theta)$ from the counterfactual input pair $(X_{i}^{\prime},X_{i}^{\prime\prime})$, $v_{1},v_{2}$ are recommendations from $\hat{R}_{i}^{\prime}$, and $rank(v,\hat{R}_{i})$ denotes the rank of $v$ in $\hat{R}_{i}$. As with SERP-K, we use $\min(\cdot,\cdot)$ to achieve symmetry.
In general, bias and fairness assessments of LLM use cases do not require satisfying all possible evaluation metrics. Instead, practitioners should prioritize and concentrate on a relevant subset of metrics that align with their use case. To demystify metric choice for these assessments, we introduce a decision framework that enables practitioners to determine suitable choices of bias and fairness evaluation metrics, drawing inspiration from the classification fairness framework proposed by Saleiro et al. (2018).
Our proposed framework is designed for use cases for which prompts can be sampled from a known population and the task is well-defined. We categorize use cases into three distinct groups based on task: 1) text generation and summarization, 2) classification, and 3) recommendation. For each category, we map use cases to a set of evaluation metrics to assess the applicable bias and fairness risks. This mapping is depicted in Figure 1, and a comprehensive list of bias and fairness evaluation metrics is contained in Table 2.
First, consider text generation and summarization. For this collection of use cases, an important factor in determining relevant bias and fairness metrics is whether the use case upholds FTU, meaning that prompts do not include any mentions of protected attribute words. If FTU is not satisfied, we recommend that practitioners include counterfactual fairness and stereotype metrics, as respectively outlined in Sections 3.1.3 and 3.1.2, in their assessments. [^26] Note that counterfactual similarity may be too strict in some contexts, particularly in use cases where textual content ought to differ across protected attribute groups (e.g. clinical contexts). For these use cases, practitioners may opt to focus only on counterfactual sentiment and omit counterfactual similarity metrics from their assessment.  [^27] Note that stereotype risk, while low, may still exist even if FTU is satisfied. Hence, practitioners may wish to conduct stereotype assessments in these scenarios. Additionally, we recommend that all text generation and summarization use cases undergo toxicity evaluation, as outlined in Section 3.1.1, regardless of whether or not FTU is upheld.
For classification use cases, we adopt a modified version of the decision framework proposed by Saleiro et al. (2018). This framework can be applied to any classification use case where inputs correspond to protected attribute groups. Following Saleiro et al. (2018), the following approach is recommended: if fairness requires that model predictions exhibit approximately equal predicted prevalence across different groups, representation fairness metrics should be used; Otherwise, error-based fairness metrics should be used. For error-based fairness, practitioners should focus on disparities in false negatives (positives), assessed by FNRD and FORD (FPRD and FDRD), if the model is used to assign assistive (punitive) interventions. [^28] In the context of fairness, if interventions are punitive, and hence can hurt individuals, it is undesirable for a model to produce false positives disproportionately for any protected attribute group. Analogously, having a model produce false negatives disproportionately for any protected attribute group is undesirable in the case of assistive interventions. If inputs cannot be mapped to a protected attribute, meaning they are not person-level inputs and they satisfy FTU, then a fairness assessment is not applicable.
Lastly, for recommendation use cases, counterfactual unfairness is a risk if FTU cannot be satisfied, as shown by Zhang et al. (2023). Note that counterfactual invariance may not be a desirable property for certain recommendation use cases. For instance, it may be preferred to recommend different products for male vs. female customers. Therefore, if counterfactual invariance is a desired property, we recommend that recommendation use cases not satisfying FTU be assessed for counterfactual unfairness in recommendations using the metrics outlined in Section 3.3. Conversely, if a recommendation use case satisfies FTU or if counterfactual invariance is not desired, then a fairness assessment is not applicable.
Figure 1: Bias and Fairness Evaluation Framework for LLM Use Cases
5 Experiments
We conduct a set of experiments using this paper’s companion Python toolkit, LangFair. [^29] We use LangFair v0.3.1 for our experiments. Given the vast number of studies that have investigated classification fairness (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023; Feldman et al., 2014; Zhang et al., 2018; Kamishima et al., 2012) and recommendation fairness (Wang et al., 2023; Li et al., 2023b; Zhang et al., 2023; Beutel et al., 2019), we focus on evaluating bias and fairness for text generation and summarization use cases. In particular, we sample from three populations of prompts and use two different LLMs, for a total of six use cases. The first two samples, each comprised of 1000 incomplete sentences, are randomly drawn from the RealToxicityPrompts (RTP) dataset (Gehman et al., 2020). [^30] In practice, we recommend practitioners adopt the sample sizes used by Wang et al. (2024b), ideally sampling 1000 prompts from their use case and generating 25 responses per prompt. If obtaining a sample size this large is not feasible, practitioners may generate a larger number of responses per prompt. One sample includes prompts with a toxicity level less than 0.2, and the other includes prompts labeled as ‘challenging’. Each of these prompts contains an incomplete sentence prepended with instructions to complete the sentence. The third sample contains 1000 conversations drawn from the DialogSum (DS) dataset (Chen et al., 2021) prepended with instructions to summarize. For the LLMs, we use gemini-1.0-pro and gpt-3.5-turbo-16k.
To select evaluation metrics, we use our decision framework depicted in Figure 1. Since we are dealing with text generation and summarization use cases, we must first determine whether our use cases satisfy FTU. Using LangFair’s CounterfactualGenerator class, we parse our three samples of prompts for gender words and find that none of the use cases satisfy FTU. We further suppose that counterfactual invariance is required for fairness in our use cases. Hence, our recommended assessments are toxicity, stereotype, and counterfactual fairness assessments. For all six use cases, we compute the full suite of toxicity, stereotype, and counterfactual metrics presented in Sections 3.1.1, 3.1.2, and 3.1.3, respectively. All results are presented in Table 3, and code snippets are contained in Appendix A.
For each use case, we generate 25 responses for each prompt and compute toxicity and stereotype metrics on the generated responses. In the toxicity assessments, we see substantial variation in toxic fraction across prompt sets for the same model and vice versa. For instance, toxic fraction is approximately 73 times higher when using gpt-3.5-turbo-16k for sentence completion with the challenging sample of RTP prompts compared to the low-toxicity sample. Upon investigation of responses with the highest toxicity scores, we find many highly offensive responses for the RTP use cases. Next, we find co-occurrence-based stereotype metrics are far more consistent across use cases, and all values are relatively low in comparison to those found in Bordia and Bowman (2019) and Liang et al. (2023). However, we find greater variation in stereotype fraction across use cases. Manual inspection of responses with highest stereotype scores reveals no cause for concern for the DS use cases, but reveals offensive content for several of the RTP use cases. [^31] LangFair provides functionality to compute response-level toxicity, stereotype, and counterfactual scores.
Lastly, we conduct a counterfactual fairness assessment for gender. Among the 1000 prompts sampled from the RTP-challenging, RTP-nontoxic, and DS datasets, 291, 189, and 306 prompts contain gender words, respectively. We subset the prompts to retain only prompts mentioning gender words, and create counterfactual input pairs (CIPs) using token-wise substitution. For each CIP, we then generate 25 responses per prompt. We find far more counterfactual variation for the use cases using gemini-1.0-pro relative to gpt-3.5-turbo-16k. Additionally, we find counterfactual responses for DS to be more similar than for either of the RTP prompt samples. This is likely due to more opportunity for creativity in sentence completion compared to dialogue summarization. After investigating response-level counterfactual scores, we find many instances with large sentiment disparities for the counterfactual responses.
6 Conclusions
In this paper, we present an actionable decision framework for selecting bias and fairness evaluation metrics for LLM use cases, introducing several new evaluation metrics as part of the framework. This work addresses two gaps in the current literature. First, to the best our knowledge, the current literature does not offer a framework for selecting bias and fairness evaluation metrics for LLM use cases. Our framework, inspired by Saleiro et al. (2018), fills this gap by incorporating use case characteristics and stakeholder values to guide the selection of evaluation metrics. Second, our framework tackles limitations of existing LLM bias and fairness evaluation approaches that rely on benchmark datasets containing predefined prompts. Instead, our approach uses actual prompts from the practitioner’s use case. By considering both prompt-risk and the assigned task of the LLM, our approach provides a more customized risk assessment for the practitioner’s specific use case. Furthermore, our proposed framework is highly practical, as all evaluation metrics are computed solely from the LLM output. To streamline implementation of the framework, all metrics can be easily computed using this paper’s companion Python toolkit, LangFair. Finally, our experiments reveal substantial variation in bias and fairness across use cases, underscoring the importance of conducting these assessments at the use case level.
7 Limitations
Despite its strengths, we note two primary limitations of this work. First, while our framework aims to encompass the vast majority of use cases for LLMs, we acknowledge that our taxonomy of use cases may not be exhaustive. Second, our framework is limited to use cases where prompts are drawn from a known population and does not cater to scenarios where the prompt population is undefined. For example, in the context of LLM chatbot applications, it’s unlikely that practitioners can dictate the prompts users input into the chatbot. Consequently, our evaluation framework is unable to account for worst-case scenarios in such use cases, where prompts could contain any text input. [^32] For more robust bias and fairness monitoring, practitioners may also consider tracking response-level scores. Using these scores, practitioners can implement response filtering or targeted human-in-the-loop practices, requiring manual review for any response that yields a concerning score. This approach offers advantages over arbitrary human-in-the-loop, where random samples of responses are manually reviewed, and exhaustive human-in-the-loop, which may be infeasible for certain large-scale use cases. [^33] Response-level toxicity and stereotype scores can be directly obtained from the respective classifiers. For counterfactual fairness, CROUGE-L, CBLEU, and CCS may be computed at the response-pair level and disparities in sentiment scores can be calculated using differences or ratios of sentiment-classifier outputs. These approaches can be implemented with the LangFair library.
Acknowledgements
We wish to thank Mohit Singh Chauhan, Blake Aber, Piero Ferrante, Xue (Crystal) Gu, Almira Pillay, Zeya Ahmad, Kee Siong Ng, Huiwen Hu, and Vasistha Singhal Vinod for their helpful suggestions as well as David Skarbrevik and Viren Bajaj for their contributions to the LangFair library.
Appendix A Code Snippets
Below, we provide code snippets from our experiments using LangFair. Note that the printed metric values are purely illustrative.
# Construct LangChain LLM
from langchain_google_vertexai import ChatVertexAI
from langchain_core.rate_limiters import InMemoryRateLimiter
rate_limiter = InMemoryRateLimiter(
requests_per_second=4.5, check_every_n_seconds=0.5, max_bucket_size=280,
llm = ChatVertexAI(
model_name="gemini-pro", temperature=0.3, rate_limiter=rate_limiter
# Generate 25 LLM responses per prompt
from langfair.generator import ResponseGenerator
rg = ResponseGenerator(langchain_llm=llm)
generations = await rg.generate_responses(prompts=prompts, count=25)
responses = generations["data"]["response"]
duplicated_prompts = generations["data"]["prompt"] # so prompts correspond to responses
# Compute toxicity metrics
import torch
from langfair.metrics.toxicity import ToxicityMetrics
device = torch.device("cuda")
tm = ToxicityMetrics(device=device)
tox_result = tm.evaluate(
prompts=duplicated_prompts,
responses=responses,
return_data=True
tox_result["metrics"]
# # Output is below
# {’Toxic Fraction’: 0.0004,
# ’Expected Maximum Toxicity’: 0.013845130120171235,
# ’Toxicity Probability’: 0.01}
# Compute stereotype metrics
from langfair.metrics.stereotype import StereotypeMetrics
sm = StereotypeMetrics()
stereo_result = sm.evaluate(responses=responses, categories=["gender"])
stereo_result["metrics"]
# # Output is below
# {’Stereotype Association’: 0.3172750176745329,
# ’Cooccurrence Bias’: 0.44766333654278373,
# ’Stereotype Fraction - gender’: 0.08}
# Check for FTU
from langfair.generator.counterfactual import CounterfactualGenerator
cg = CounterfactualGenerator(langchain_llm=llm)
ftu_result = cg.check_ftu(
prompts=prompts,
attribute="gender",
subset_prompts=True
pd.DataFrame(ftu_result["data"])
# Generate counterfactual responses
cf_generations = await cg.generate_responses(
prompts=prompts, attribute="gender", count=25
male_responses = cf_generations["data"]["male_response"]
female_responses = cf_generations["data"]["female_response"]
# Compute counterfactual metrics
from langfair.metrics.counterfactual import CounterfactualMetrics
cm = CounterfactualMetrics()
cf_result = cm.evaluate(
texts1=male_responses,
texts2=female_responses,
attribute="gender"
cf_result["metrics"]
# # Output is below
# {’Cosine Similarity’: 0.8318708,
# ’RougeL Similarity’: 0.5195852482361165,
# ’Bleu Similarity’: 0.3278433712872481,
# ’Sentiment Bias’: 0.0009947145187601957}
Appendix B Supplemental Tables
Table 1: Taxonomy of LLM Use Cases and Associated Bias/Fairness Risks
Use Case Category Description Examples Bias/Fairness Risk
Table 2: Glossary of Bias and Fairness Evaluation Metrics